# Shifted Interpolation for Differential Privacy

### v1

@@ 1 Introduction
In words, DP measures how distinguishable the output of a (randomized) algorithm is when run on two adjacent datasets, i.e., two datasets that differ in only one individual record. There are several ways to measure distinguishability—leading to many relaxations of DP, e.g., [BS16, Mir17, DRS22]. Different DP notions lead to different privacy analyses, and a long line of work has sought to prove sharp privacy bounds for noisy gradient descent and its variants [BST14, ${}^{+}$16, FMTT18, CYS21, YS22, AT22, ABT24].
@@ 1.1 Contribution
Notably, this yields the first exact privacy analysis in the foundational setting of strongly convex losses. To our knowledge, there is no other setting where exact privacy analyses are known for $t>1$, except for the setting of convex quadratic losses which is analytically trivial because all iterates are explicit Gaussians. [^2] The standard analysis approach based on the composition theorem is nearly tight for small numbers of iterations $t$, but as mentioned above, yields an arbitrarily loose bound (in fact vacuous) for convex losses as $t\to\infty$.
@@
Since our improved privacy guarantees are for $f$-DP (Figure 1, left), lossless conversions immediately imply improved guarantees for other notions of privacy like Rényi DP and $(\varepsilon,\delta)$-DP (Figure 1, right). For example, for the strongly convex setting, our exact bound improves over previous results by a factor of $2$ in Rényi DP, and thus by even more in $(\varepsilon,\delta)$-DP due to the intrinsic lossiness of Rényi DP that we overcome by directly analyzing in $f$-DP. In practice, improving the privacy by a factor of two enables training for twice as long or with half the noise, while satisfying the same privacy budget. Although this paper’s focus is the theoretical methodology, preliminary numerics in §4.4 corroborate that our improved privacy guarantees can be helpful in practice.
Figure 1: Left: improved $f$-DP vs. the standard composition analysis. Right: improved $(\varepsilon,\delta)$-DP by losslessly converting from $f$-DP. Our privacy bound is optimal in all parameters, here for NoisyGD on strongly convex losses; see §D for the parameter choices and other settings. Our $f$-DP analysis also implies optimal bounds for the Rényi DP framework, but $f$-DP is strictly better since it captures all aspects of the privacy leakage, whereas Rényi DP is intrinsically lossy.
@@ 1.2 Techniques
A key motivation behind the construction of our auxiliary sequence is that it demystifies the popular privacy amplification by iteration analysis [FMTT18], which has been used in many contexts, and in particular was recently shown to give convergent Rényi DP bounds [AT22, ABT24]. Those arguments rely on shifted divergences, which combine Rényi divergence and Wasserstein distance, and it was an open question whether this ad-hoc potential function could be simplified. Our shifted interpolated process answers this: its iterates coincide with the optimal “shifts” in the shifted divergence argument, which allows us to disentangle the Rényi and Wasserstein components of the shifted divergence argument; details in §B. Crucially, this disentanglement enables generalizations beyond divergence-based relaxations of DP, to $f$-DP. [^3] Naïvely extending the “shifted divergence” argument to “shifted tradeoffs” runs into subtle but fundamental issues since tradeoff functions do not enjoy key properties that information divergences do. Details in §B.
@@ 1.3 Outline
§2 recalls relevant preliminaries from differential privacy and convex optimization. §3 introduces the technique of shifted interpolation. §4 uses this to establish improved privacy bounds for noisy gradient descent and its variants for the foundational settings of convex and strongly convex losses. §5 describes how, as immediate corollaries of these convergent privacy bounds, taking an appropriate limit recovers and generalizes recent results on the $f$-DP of the exponential mechanism. §6 discusses future directions motivated by our results. For brevity, various helper lemmas, proof details, and additional numerical experiments are deferred to the Appendix. Code reproducing our numerics can be found here: https://github.com/jinhobok/shifted_interpolation_dp.
@@ 2.1 Differential privacy
See §A.2 for further details on tradeoff functions. Gaussian tradeoff functions are a particularly useful family. These are central to our analysis due to the Gaussian noise in noisy (stochastic) gradient descent.
@@
The composition of two tradeoff functions $f=T(P,Q)$ and $g=T(P^{\prime},Q^{\prime})$ is defined as $f\otimes g=T(P\times P^{\prime},Q\times Q^{\prime})$. The $n$-fold composition of $f$ with itself is denoted $f^{\otimes n}$.
@@ 2.3 Private optimization algorithms
A family of loss functions $\mathcal{F}$ has gradient sensitivity $L$ if
$\sup_{f,g\in\mathcal{F}}\lVert{\nabla f-\nabla g}\rVert_{\infty}\leq L\,.$
@@ 3 Shifted interpolation for $f$-DP
Here we explain the key conceptual ideas enabling our convergent $f$-DP bounds (see §1.2 for a high-level discussion). To preserve the flow of ideas we defer proofs to the Appendix. Below, in §3.1 we first recall the standard $f$-DP analysis based on the composition theorem and why it yields divergent bounds. Then in §3.2 we describe our technique of shifted interpolated processes and how this enables convergent $f$-DP bounds.
@@ 3.1 Previous (divergent) $f$-DP bounds, via composition
Suppose $\lVert{\phi-\phi^{\prime}}\rVert_{\infty}\leq s$. Then
@@ 3.2 Convergent $f$-DP bounds, via shifted interpolation
Intuitively, this argument replaces the divergent $\sqrt{t}$ dependence of prior $f$-DP bounds with something scaling in $t-\tau$. Here $\tau$ is an analysis parameter that we can optimize based on the following intuitive tradeoff: larger $\tau$ enables unrolling less, whereas smaller $\tau$ gives the auxiliary process $\{\widetilde{X}_{k}\}$ more time to interpolate between $X_{\tau}^{\prime}$ and $X_{t}$ while yielding smaller penalties $a_{k}$ for unrolling at each iteration.
@@ 3.2.2 Geometrically aware composition
Suppose that $\|\phi-\phi^{\prime}\|_{\infty}\leq s$ and that $\phi,\phi^{\prime}$ are $c$-Lipschitz. Then for any $\lambda\geq 0$ and any random variable $\widetilde{X}$ satisfying $\lVert{X-\widetilde{X}}\rVert\leq z$,
@@ 3.2.3 Convergent $f$-DP bounds
Consider the stochastic processes $\{X_{k}\},\{X_{k}^{\prime}\},\{\widetilde{X}_{k}\}$ defined in (3.1), (3.2), (3.5), with $\lambda_{t}=1$. Suppose that $\|\phi-\phi^{\prime}\|_{\infty}\leq s$ and that $\phi,\phi^{\prime}$ are $c$-Lipschitz. For any sequence $\{z_{k}\}$ such that $\lVert{X_{k}-\widetilde{X}_{k}}\rVert\leq z_{k}$,
@@ 4.1 Noisy gradient descent
Theorem 4.2 is exactly tight in all parameters, and improves over the composition-based analysis (Theorem 4.1) for all $t>1$. See Figure 1Theorem 4.3 is tight up to a constant factor [AT22], and for $t>\frac{4Dn}{\eta L}$ it dominates Theorem 4.1 since its convergent nature outweighs the slightly suboptimal constant.
@@ 4.3 Noisy stochastic gradient descent
Compared to NoisyGD, the privacy leakage in NoisySGD only occurs when the index $i^{*}$ is in the sampled batch. This phenomenon is known as privacy amplification by subsampling [${}^{+}$11], which is formulated in $f$-DP as follows [DRS22, Definition 6].
@@
[…]
@@
Both theorems give convergent privacy by taking $t-\tau$ constant as $t\to\infty$. In contrast, Theorem 4.8 is convergent in the regime $t=O(n^{2}/b^{2})$ (by CLT), but yields a vacuous privacy as $t\to\infty$ for fixed $b/n$ [DRS22]. We remark that for finite but large $t$, one can set $t-\tau$ to be sufficiently large and apply CLT (A.5) to approximate the composition of $C_{p}(G(\cdot))$; see A.11. We also remark that by choosing $t-\tau=\Theta(\tfrac{Dn}{\eta L})$, Theorem 4.10 recovers the asymptotically tight Rényi DP bound of [AT22].
@@ 4.4 Numerical example
As a proof of concept, here we consider regularized logistic regression on MNIST [LCB10]. We compare our results with the state-of-the-art Rényi DP bounds, and existing $f$-DP bounds (based on the composition theorem) which we denote as GDP Composition. For a fair comparison, we use the same algorithm NoisyCGD, with all parameters unchanged, and only focus on the privacy accounting; we focus on NoisyCGD because it is close to standard private optimization implementations, e.g., Pytorch [${}^{+}$19] and TensorFlow [${}^{+}$16]. Indeed, standard implementations often cycle through batches in a permuted order every epoch, somewhat similar to how NoisyCGD cycles through batches in a fixed order.
@@ 5.1 Strongly log-concave targets
It is well-known that under these assumptions, LMC converges to its stationary distribution in total variation (TV) distance as $t\to\infty$, and the stationary distribution converges to the exponential mechanism as $\eta\to 0$, see e.g., [Che23]. By A.10, tradeoff functions converge under TV. ∎
@@ 6 Discussion
One natural direction is whether convergent $f$-DP bounds can be shown in more general settings, e.g., (structured) non-convex landscapes, heteroscedastic or correlated noise [${}^{+}$23], adaptive first-order algorithms, or second-order algorithms [GHST23].
A technical question is whether one can relax the $W_{\infty}$ bounds between our shifted interpolated process $\{\widetilde{X}_{k}\}$ and the target process $\{X_{k}\}$, and if this can enable tighter analyses of stochastic algorithms. While $W_{\infty}$ has traditionally been used for privacy amplification by iteration [FMTT18][AC23a] recently showed that some of this analysis extends to the Orlicz–Wasserstein distance, and is even necessary in some applications.
@@ Acknowledgements.
We thank Sinho Chewi, Kunal Talwar, and Jiayuan Ye for insightful conversations about the literature. We also thank Jiayuan Ye for sharing helpful code for numerical experiments.
@@ A.1 Rényi DP
Conversion from RDP to $(\varepsilon,\delta)$-DP is inherently lossy and there are many proposed formulae for this. Since the RDP bounds mentioned in this text are of the form $(\alpha,\rho\alpha)$-RDP for all $\alpha>1$, given $\rho$ and a fixed level of $\delta$ the corresponding converted value of $\varepsilon=\varepsilon(\alpha,\rho,\delta)$ can be found by optimizing over $\alpha$. Also, in addition to RDP, results on zero concentrated DP [BS16] can be applied. Throughout, we calculate the minimum $\varepsilon$ (aka the best bound) among the following formulae: [BS16, Lemma 3.5], [Mir17, Proposition 3], [${}^{+}$20, Theorem 20], and [${}^{+}$21, Lemma 1].
@@ A.3 Convergence of tradeoff functions
[…]
@@ Appendix B Disentangling the shift in shifted divergences
For (2), the disentangling we achieve in (1) appears essential. The naïve approach of directly extending the shifted divergence argument to “shifted tradeoffs” $T^{(z)}(P,Q)=\sup_{P^{\prime}:W_{\infty}(P,P^{\prime})\leq z}(P^{\prime},Q)$ runs into several subtle technical issues. For example, the argument appears to require the existence of an optimal shift $P^{\prime}$. For the shifted Rényi argument, it suffices to find a nearly-optimal shift $D_{\alpha}^{(z)}(P\mathbin{\|}Q)=\inf_{P^{\prime}:W_{\infty}(P,P^{\prime})\leq z}D_{\alpha}(P^{\prime}\mathbin{\|}Q)$, and moreover have the shift be nearly-optimal for a given Rényi parameter $\alpha$ but perhaps not uniformly so over all $\alpha$. Due to the more involved calculus of tradeoff functions, these issues become subtle but important problems, and have led others to state the problem of privacy amplification by iteration in $f$-DP as open, e.g., [${}^{+}$23]. Although the general problem of finding an optimal shift for general tradeoff functions remains open, the answer to (1)—our shifted interpolated process—explicitly constructs an optimal shift for the tradeoff functions specifically needed to analyze two contractive noisy iterations.

### v2

@@ 1 Introduction
In words, DP measures how distinguishable the output of a (randomized) algorithm is when run on two adjacent datasets, i.e., two datasets that differ in only one individual record. There are several ways to measure distinguishability—leading to many relaxations of DP, e.g., [BS16, Mir17, DRS22]. Different DP notions lead to different privacy analyses, and a long line of work has sought to prove sharp privacy bounds for noisy gradient descent and its variants [BST14, ACG+16, FMTT18, CYS21, YS22, AT22, ABT24].
@@ 1.1 Contribution
Notably, this yields the first exact privacy analysis in the foundational setting of strongly convex losses. To our knowledge, there is no other setting where exact privacy analyses are known for any $t>1$, except for the setting of convex quadratic losses which is analytically trivial because all iterates are explicit Gaussians. [^2] The standard analysis approach based on the composition theorem is nearly tight for small numbers of iterations $t$, but as mentioned above, yields an arbitrarily loose bound (in fact vacuous) for convex losses as $t\to\infty$.
@@
Since our improved privacy guarantees are for $f$-DP (Figure 1, left), lossless conversions immediately imply improved guarantees for other notions of privacy like Rényi DP and $(\varepsilon,\delta)$-DP (Figure 1, right). For example, for the strongly convex setting, our exact bound improves over previous results by a factor of $2$ in Rényi DP, and thus by even more in $(\varepsilon,\delta)$-DP due to the intrinsic lossiness of Rényi DP that we overcome by directly analyzing in $f$-DP. In practice, improving the privacy by a factor of two enables training with half the noise, while satisfying the same privacy budget. Although this paper’s focus is the theoretical methodology, preliminary numerics in §4.4 corroborate that our improved privacy guarantees can be helpful in practice.
Figure 1: Left: improved $f$-DP versus the standard composition analysis. Right: improved $(\varepsilon,\delta)$-DP by losslessly converting from $f$-DP. Our privacy bound is optimal in all parameters, here for NoisyGD on strongly convex losses; see §D for the parameter choices and other settings. Our $f$-DP analysis also implies optimal bounds for the Rényi DP framework (previously unknown), but $f$-DP is strictly better since it captures all aspects of the privacy leakage, whereas Rényi DP is intrinsically lossy.
@@ 1.2 Techniques
A key motivation behind the construction of our auxiliary sequence is that it demystifies the popular privacy amplification by iteration analysis [FMTT18], which has been used in many contexts, and in particular was recently shown to give convergent Rényi DP bounds [AT22, ABT24]. Those arguments rely on shifted divergences, which combine Rényi divergence and Wasserstein distance, and it was an open question whether this ad-hoc potential function could be simplified. Our shifted interpolated process answers this: its iterates coincide with the optimal “shifts” in the shifted divergence argument, which allows us to disentangle the Rényi and Wasserstein components of the shifted divergence argument; details in §B. Crucially, this disentanglement enables generalizations beyond divergence-based relaxations of DP, to $f$-DP. [^3] Naïvely extending the “shifted divergence” argument to “shifted tradeoffs” runs into subtle but fundamental issues since tradeoff functions do not enjoy key properties that divergences do. Details in §B.
@@ 1.3 Outline
§2 recalls relevant preliminaries from differential privacy and convex optimization. §3 introduces the technique of shifted interpolation. §4 uses this to establish improved privacy bounds for noisy gradient descent and its variants in a number of settings. §5 describes how, as immediate corollaries of these convergent privacy bounds, taking an appropriate limit recovers and generalizes recent results on the $f$-DP of the exponential mechanism. §6 discusses future directions motivated by our results. For brevity, various helper lemmas, proof details, and additional numerical experiments are deferred to the Appendix. Code reproducing our numerics can be found here: https://github.com/jinhobok/shifted_interpolation_dp.
@@ 2.1 Differential privacy
See §A.2 for further details on tradeoff functions. Gaussian tradeoff functions are a particularly useful family, providing a notion of Gaussian DP (GDP) parametrized by a single scalar. These are central to our analysis due to the Gaussian noise in noisy (stochastic) gradient descent.
@@
The composition of two tradeoff functions $f=T(P,Q)$ and $g=T(P^{\prime},Q^{\prime})$ is defined as $f\otimes g=T(P\times P^{\prime},Q\times Q^{\prime})$. The $n$-fold composition of $f$ with itself is denoted as $f^{\otimes n}$.
@@ 2.3 Private optimization algorithms
A family of loss functions $\mathcal{F}$ (defined on $\mathcal{X}$) has gradient sensitivity $L$ if
$\sup_{f,g\in\mathcal{F},x\in\mathcal{X}}\lVert{\nabla f(x)-\nabla g(x)}\rVert\leq L\,.$
@@ 3 Shifted interpolation for $f$-DP
Here we explain the key conceptual ideas enabling our convergent $f$-DP bounds (see §1.2 for a high-level discussion). To preserve the flow of ideas we defer proofs to §C.1. Below, in §3.1 we first recall the standard $f$-DP analysis based on the composition theorem and why it yields divergent bounds. Then in §3.2 we describe our technique of shifted interpolated processes and how this enables convergent $f$-DP bounds.
@@ 3.1 Previous (divergent) $f$-DP bounds, via composition
Suppose $\lVert{\phi(x)-\phi^{\prime}(x)}\rVert\leq s$ for all $x$. Then
@@ 3.2 Convergent $f$-DP bounds, via shifted interpolation
Intuitively, this argument replaces the divergent $\sqrt{t}$ dependence of prior $f$-DP bounds with something scaling in $t-\tau$. Here $\tau$ is an analysis parameter that we can optimize based on the following intuitive tradeoff: larger $\tau$ enables unrolling less, whereas smaller $\tau$ gives the auxiliary process $\{\widetilde{X}_{k}\}$ more time to interpolate between $X_{\tau}^{\prime}$ and $X_{t}$ which leads to smaller penalties $a_{k}$ for unrolling at each iteration.
@@ 3.2.2 Geometrically aware composition
Suppose that $\lVert{\phi(x)-\phi^{\prime}(x)}\rVert\leq s$ for all $x$ and that $\phi,\phi^{\prime}$ are $c$-Lipschitz. Then for any $\lambda\geq 0$ and any random variable $\widetilde{X}$ satisfying $\lVert{X-\widetilde{X}}\rVert\leq z$,
@@ 3.2.3 Convergent $f$-DP bounds
Consider the stochastic processes $\{X_{k}\},\{X_{k}^{\prime}\},\{\widetilde{X}_{k}\}$ defined in (3.1), (3.2), (3.5), with $\lambda_{t}=1$. Suppose that $\lVert{\phi(x)-\phi^{\prime}(x)}\rVert\leq s$ for all $x$ and that $\phi,\phi^{\prime}$ are $c$-Lipschitz. For any sequence $\{z_{k}\}$ such that $\lVert{X_{k}-\widetilde{X}_{k}}\rVert\leq z_{k}$,
@@ 4.1 Noisy gradient descent
Theorem 4.2 is exactly tight in all parameters, and improves over the composition-based analysis (Theorem 4.1) for all $t>1$. See Figure 1Theorem 4.3 is tight up to a constant factor (see Theorem C.16), and for $t>\frac{4Dn}{\eta L}$ it dominates Theorem 4.1 since its convergent nature outweighs the slightly suboptimal constant.
@@ 4.3 Noisy stochastic gradient descent
Compared to NoisyGD, the privacy leakage in NoisySGD only occurs when the index $i^{*}$ is in the sampled batch. This phenomenon is known as privacy amplification by subsampling [KLN+11], which is formulated in $f$-DP as follows [DRS22, Definition 6].
@@
For tradeoff function $f$ and $p\in[0,1]$, define $f_{p}=pf+(1-p)\text{Id}$. The subsampling operator $C_{p}$ (with respect to $f$) is defined as $C_{p}(f)=\min\{f_{p},(f_{p})^{-1}\}^{**}$ where -1 denotes the (left-continuous) inverse and denotes the convex conjugate. Equivalently, $C_{p}(f)$ is the pointwise largest tradeoff function $g$ such that $g\leq f_{p}$ and $g\leq(f_{p})^{-1}$.
@@
Both theorems give convergent privacy by taking $t-\tau$ constant as $t\to\infty$. In contrast, Theorem 4.8 is convergent in the regime $t=O(n^{2}/b^{2})$ (by CLT), but yields a vacuous privacy as $t\to\infty$ for fixed $b/n$ [DRS22]. We remark that for finite but large $t$, one can set $t-\tau$ to be sufficiently large and apply CLT (A.5) to approximate the composition of $C_{p}(G(\cdot))$; see A.11 and § C.4.3. We also remark that by choosing $t-\tau=\Theta(\tfrac{Dn}{\eta L})$, Theorem 4.10 recovers the asymptotically tight Rényi DP bound of [AT22].
@@ 4.4 Numerical example
As a proof of concept, here we consider regularized logistic regression on MNIST [LCB10]. We compare our results with the state-of-the-art Rényi DP bounds, and existing $f$-DP bounds (based on the composition theorem) which we denote as GDP Composition. For a fair comparison, we use the same algorithm NoisyCGD, with all parameters unchanged, and only focus on the privacy accounting; we focus on NoisyCGD because it is close to standard private optimization implementations, e.g., Pytorch [PGM+19] and TensorFlow [ABC+16]. Indeed, standard implementations often cycle through batches in a permuted order every epoch, somewhat similar to how NoisyCGD cycles through batches in a fixed order.
@@ 5.1 Strongly log-concave targets
It is well-known that under these assumptions, LMC converges to its stationary distribution in total variation (TV) distance as $t\to\infty$, and the stationary distribution converges to the exponential mechanism as $\eta\to 0$, see e.g., [Che23]. By A.10, tradeoff functions converge under convergence in TV. ∎
@@ 6 Discussion
One natural direction is whether convergent $f$-DP bounds can be shown in more general settings, e.g., (structured) non-convex landscapes, heteroscedastic or correlated noises [CCDP+23], adaptive first-order algorithms, or second-order algorithms [GHST23].
A technical question is whether one can relax the $W_{\infty}$ bounds between our shifted interpolated process $\{\widetilde{X}_{k}\}$ and the target process $\{X_{k}\}$, and if this can enable tighter analyses of stochastic algorithms. While $W_{\infty}$ has traditionally been used for privacy amplification by iteration [FMTT18][AC23a] recently showed that some of this analysis extends to the Orlicz–Wasserstein distance, which is even necessary in some applications.
@@ Acknowledgements.
We thank Sinho Chewi, Kunal Talwar, and Jiayuan Ye for insightful conversations about the literature and the anonymous reviewers for helpful comments. We also thank Jiayuan Ye for sharing helpful code for numerical experiments.
@@ A.1 Rényi DP
Conversion from RDP to $(\varepsilon,\delta)$-DP is inherently lossy and there are many proposed formulae for this. Since the RDP bounds mentioned in this text are of the form $(\alpha,\rho\alpha)$-RDP for all $\alpha>1$, given $\rho$ and a fixed level of $\delta$ the corresponding converted value of $\varepsilon=\varepsilon(\alpha,\rho,\delta)$ can be found by optimizing over $\alpha$. Also, in addition to RDP, results on zero concentrated DP [BS16] can be applied. Throughout, we calculate the minimum $\varepsilon$ (aka the best bound) among the following formulae: [BS16, Lemma 3.5], [Mir17, Proposition 3], [BBG+20, Theorem 20], and [ALC+21, Lemma 1].
@@ A.3 Convergence of tradeoff functions
The restriction $\alpha^{\prime}\in(0,1]$ is necessary. For example, if $P=\delta_{0}$, $Q=\tfrac{1}{2}\delta_{0}+\tfrac{1}{2}\delta_{1}$, and $P_{n}=(1-\tfrac{1}{n})\delta_{0}+\tfrac{1}{n}\delta_{1}$ (here, $\delta_{x}$ denotes the Dirac measure at $x$ and $pP+(1-p)Q$ denotes the mixture of $(P,Q)$ with mixing rate $(p,1-p)$), then $\text{TV}(P_{n},P)\to 0$ implies $T(P_{n},P)\to\text{Id}$ and $T(P,Q)(\alpha)=\tfrac{1}{2}(1-\alpha)$, yet
@@ Appendix B Disentangling the shift in shifted divergences
For (2), the disentangling we achieve in (1) appears essential. The naïve approach of directly extending the shifted divergence argument to “shifted tradeoffs” $T^{(z)}(P,Q)=\sup_{P^{\prime}:W_{\infty}(P,P^{\prime})\leq z}(P^{\prime},Q)$ runs into several subtle technical issues. For example, the argument appears to require the existence of an optimal shift $P^{\prime}$. For the shifted Rényi argument, it suffices to find a nearly-optimal shift $D_{\alpha}^{(z)}(P\mathbin{\|}Q)=\inf_{P^{\prime}:W_{\infty}(P,P^{\prime})\leq z}D_{\alpha}(P^{\prime}\mathbin{\|}Q)$, and moreover have the shift be nearly-optimal for a given Rényi parameter $\alpha$ but perhaps not uniformly so over all $\alpha$. Due to the more involved calculus of tradeoff functions, these issues become subtle but important problems, and have led others to state the problem of privacy amplification by iteration in $f$-DP as open, e.g., [WSY+23]. Although the general problem of finding an optimal shift for general tradeoff functions remains open, the answer to (1)—our shifted interpolated process—explicitly constructs an optimal shift for the tradeoff functions specifically needed to analyze two contractive noisy iterations.